OpenStudy (anonymous):

could someone tell me the steps to this problem? 28abc/60ac

OpenStudy (saifoo.khan):

\[\frac{28abc}{60ac}\] \[\frac{28\cancel ab\cancel c}{60\cancel a\cancel c}\] \[\frac{28b}{60}\]

OpenStudy (saifoo.khan):

Now we have to simplify 28/60